OpenStudy (anonymous):

Find the point on the curve r(t)=<2cos(t),2sin(t),e^t>, 0

OpenStudy (anonymous):

\[<2\cos t,2\sin t, e ^{t}>,\\ 0 <= t <= \pi \\ \sqrt{3}x+y=1\]